# Benchmarking Deep Learning-Based Object Detection Models on Feature Deficient Astrophotography Imagery Dataset

### v1

@@ 3.1 Training hardware
PP-YOLOE+ PaddlePaddle Light 100
@@ 3.4 PPYoloE+
PP-YOLOE+ [17] employed a CSPResNet backbone and CSPPAN neck, trained for 100 epochs with a batch size of 8 in 32-bit precision. The learning rate schedule included a 5-epoch linear warmup followed by cosine decay from a base of 0.001. SGD was used with momentum 0.9 and L2 regularization of 0.0005. Data augmentations included geometric distortions, cropping, flipping, and multi-scale resizing. Synchronized BatchNorm and EMA (decay 0.9998) were also applied.
@@ 4 Observations
Comparing the results from our models with their COCO benchmarks reveals a marked deterioration in performance on our feature-deficient dataset for most of the models. For instance, PP-Yoloe+ exhibits a 15.9 percentage point decline in Average Precision (AP), and YOLOv12 falls from 55.2 mAP to 38.9 mAP.

### v2

@@ 3.1 Training hardware
$P$-YOLOE+ PaddlePaddle Light 100
@@ 3.4 PPYoloE+
$P$-YOLOE+ [17] employed a CSPResNet backbone and CSPPAN neck, trained for 100 epochs with a batch size of 8 in 32-bit precision. The learning rate schedule included a 5-epoch linear warmup followed by cosine decay from a base of 0.001. SGD was used with momentum 0.9 and L2 regularization of 0.0005. Data augmentations included geometric distortions, cropping, flipping, and multi-scale resizing. Synchronized BatchNorm and EMA (decay 0.9998) were also applied.
@@ 4 Observations
Comparing the results from our models with their COCO benchmarks reveals a marked deterioration in performance on our feature-deficient dataset for most of the models. For instance, $P$-Yoloe+ exhibits a 15.9 percentage point decline in Average Precision (AP), and YOLOv12 falls from 55.2 mAP to 38.9 mAP.